# Degenerate scalar scenario of two Higgs doublet model with a complex singlet scalar

### v1

@@ VI Summary and discussion
In this paper, we have investigated the compatibility of the degenerate scalar scenario and strong first-order EWPT in the 2HDMS, which contains two Higgs doublets and one singlet scalar. In this model, the imaginary part of $S$ behaves as a DM $\chi$, and CP-even scalars mix to form three Higgs particles $H_{i}~{}(i=1,2,3)$. The scattering of DM and quarks is mediated by $H_{i}$, and we have found that this scattering is suppressed in the region where Higgs masses are degenerate, as shown in previous studies on pNG DM Abe et al. (2021). The suppression of SU(2)L doublet-singlet mixing terms $\delta_{1}$ and $\delta_{2}$, which are related to the orthogonality of the mixing matrix that transforms the current eigenstates to the mass eigenstates of the CP-even scalars, is important for the degenerate scalar scenario.
@@ Acknowledgements.
We are grateful to Ayana Okuma and Chiaki Nose for their valuable discussions. The work of GCC is supported by JSPS KAKENHI Grant No. 22K03616.
@@ Appendix A Input and output parameters
The following shows the relationship between the input and output parameters. However, $m_{3}^{2}$ and $a_{1}$ remain as input parameters. The index $i$ runs 1 to 3, and the sum symbol $\quantity(\sum_{i})$ is omitted.
@@
$\lambda_{1}$ $=\frac{1}{v_{1}^{2}}\left(O_{1i}^{2}m_{H_{i}}^{2}-\frac{m_{3}^{2}v_{2}}{v_{1}}\right),$ (92)
$\lambda_{2}$ $=\frac{1}{v_{2}^{2}}\left(O_{2i}^{2}m_{H_{i}}^{2}-\frac{m_{3}^{2}v_{1}}{v_{2}}\right),$ (93)
$\lambda_{3}$ $=\frac{1}{v_{1}v_{2}}\left(O_{1i}O_{2i}m_{H_{i}}^{2}+m_{3}^{2}\right)-\lambda_{4}-\lambda_{5},$ (94)
@@
$\delta_{1}$ $=\frac{2}{v_{1}v_{s}}\left(O_{1i}O_{3i}m_{H_{i}}^{2}\right),$ (97)
$\delta_{2}$ $=\frac{2}{v_{2}v_{s}}\left(O_{2i}O_{3i}m_{H_{i}}^{2}\right),$ (98)
$d_{2}$ $=\frac{2}{v_{s}^{2}}\left(\frac{\sqrt{2}a_{1}}{v_{s}}+O_{3i}^{2}m_{H_{i}}^{2}\right),$ (99)

### v2

@@ VI Summary and discussion
In this paper, we have investigated the compatibility of the degenerate scalar scenario and strong first-order EWPT in the 2HDMS, which contains two Higgs doublets and one singlet scalar. In this model, the imaginary part of $S$ behaves as a DM $\chi$, and CP-even scalars mix to form three Higgs particles $H_{i}~{}(i=1,2,3)$. The scattering of DM and quarks is mediated by $H_{i}$, and we have found that this scattering is suppressed in the region where Higgs masses are degenerate, as shown in previous studies on pNG DM Abe et al. (2021). The suppression of SU(2)L doublet-singlet mixing terms $\delta_{1}$ and $\delta_{2}$, which are related to the orthogonality of the mixing matrix that transforms the gauge eigenstates to the mass eigenstates of the CP-even scalars, is important for the degenerate scalar scenario.
@@ Acknowledgements.
We are grateful to Ayana Okuma and Chiaki Nose for their valuable discussions. The work of GCC is supported by JSPS KAKENHI Grant No. 22K03616. The part of the work of CI is supported by the National Natural Science Foundation of China (NNSFC) Grant No. 12205387, No.12475111.
@@ Appendix A Input and output parameters
The following shows the relationship between the input and output parameters. However, $m_{3}^{2}$ and $a_{1}$ remain as input parameters. The index $i$ runs from 1 to 3, and the sum symbol $\quantity(\sum_{i})$ is omitted.
@@
$\lambda_{1}$ $=\frac{1}{v_{1}^{2}}\left(\sum_{i=1}^{3}O_{1i}^{2}m_{H_{i}}^{2}-\frac{m_{3}^{2}v_{2}}{v_{1}}\right),$ (114)
$\lambda_{2}$ $=\frac{1}{v_{2}^{2}}\left(\sum_{i=1}^{3}O_{2i}^{2}m_{H_{i}}^{2}-\frac{m_{3}^{2}v_{1}}{v_{2}}\right),$ (115)
$\lambda_{3}$ $=\frac{1}{v_{1}v_{2}}\left(\sum_{i=1}^{3}O_{1i}O_{2i}m_{H_{i}}^{2}+m_{3}^{2}\right)-\lambda_{4}-\lambda_{5},$ (116)
@@
$\delta_{1}$ $=\frac{2}{v_{1}v_{S}}\sum_{i=1}^{3}O_{1i}O_{3i}m_{H_{i}}^{2},$ (119)
$\delta_{2}$ $=\frac{2}{v_{2}v_{S}}\sum_{i=1}^{3}O_{2i}O_{3i}m_{H_{i}}^{2},$ (120)
$d_{2}$ $=\frac{2}{v_{S}^{2}}\left(\frac{\sqrt{2}a_{1}}{v_{S}}+\sum_{i=1}^{3}O_{3i}^{2}m_{H_{i}}^{2}\right),$ (121)
@@
Appendix B Trilinear and quartic couplings in the pNG-DM model
We collect trilinear and bilinear couplings of the potential (1) in the gauge eigenstate, which appear in Sec. III.
$C_{\chi\chi h_{1}}$ $=\frac{\delta_{1}}{4}v_{1},\quad C_{\chi\chi h_{2}}=\frac{\delta_{2}}{4}v_{2},\quad C_{\chi\chi s}=\frac{d_{2}}{4}v_{S},$
$C_{h_{1}s}$ $=\frac{\delta_{2}}{4}v_{1}v_{S},\quad C_{h_{2}s}=\frac{\delta_{2}}{2}v_{2}v_{S},\quad C_{ss}=\frac{b_{1}}{2}+\frac{b_{2}}{2}+\frac{\delta_{1}}{8}v_{1}^{2}+\frac{\delta_{1}}{8}v_{2}^{2}+\frac{3d_{2}}{8}v_{S}^{2}.$ (124)